# Early stopping for conjugate gradients in statistical inverse problems

### v1

@@ 9.4 Proofs for Section 5
We prove the claim analogously to Johnstone [15, Proposition 4.23]. Without loss of generality, assume that $v_{1},\dots,v_{D}$ are the first $D$ canonical unit vectors. Let $\Theta^{m}(\eta)$, $m=1,\dots,D$, for $\eta\geqslant 0$ denote the set of $f\in{\mathbb{R}}^{D}$ that satisfy $\lvert f_{i}\rvert\leqslant\eta$ for $i=1,\dots,m$ and $f_{i}=0$ for $i=m+1,\dots,D$. Define the loss
@@
Hence, for $\Theta^{m}(\delta m^{p})$ to be contained in the class of signals $f$ satisfying $\lVert f\rVert_{\mu}\leqslant R$, it suffices that $c_{A}^{-4\mu}\delta^{2}m^{4\mu p+2p+1}\leqslant R^{2}$. Let $m_{1}$ be the largest integer such that $c_{A}^{-4\mu}m_{1}^{4\mu p+2p+1}\leqslant R^{2}/\delta^{2}$ and $m_{0}\in{\mathbb{R}}$ the solution to $c_{A}^{-4\mu}m_{0}^{4\mu p+2p+1}=R^{2}/\delta^{2}$. If $m_{0}\geqslant 4$, then $m_{1}/m_{0}\geqslant 1/2$, which can be shown by contradiction. For $\delta/R\leqslant d_{\mu,p,c_{A}}$, we conclude
@@ 9.5 Proofs for Section 7
    If $\lvert r_{\tau}^{\prime}(0)\rvert<\lambda_{D}^{-2}$, then by Lemma 6.5
@@
    with probability at least $1-e^{-z}$ and we conclude by Corollary 7.3.
    If $\lvert r_{\tau}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, then we obtain the rough bound
@@
    using the prediction error bound from Proposition 5.2. Moreover, we then have by Lemma 4.7
    $S_{\tau,\lambda}\geqslant\lVert\mathsf{(1-{\exp(-\lambda_{D}^{-2}x)})^{1/2}}\xi\rVert^{2}\geqslant(1-e^{-1})\lVert\xi\rVert^{2}.$
    The lower $\chi^{2}$-concentration bound of Laurent and Massart [16, Lemma 1] gives with probability at least $1-e^{-z}$
    $\delta^{-2}\lVert\xi\rVert^{2}\geqslant D-2\sqrt{D}\sqrt{z},\quad\text{i.e.}\quad(\sqrt{D}-\sqrt{z})^{2}\leqslant\delta^{-2}\lVert\xi\rVert^{2}+z.$
    This shows by PSD($p$,$c_{A}$,$C_{A}$)
    $\lambda_{D}^{-2}\lesssim_{c_{A}}D^{2p}\leqslant(\delta^{-1}\lVert\xi\rVert+2\sqrt{z})^{4p}\lesssim_{p}(\delta^{-2}S_{\tau,\lambda})^{2p}+z^{2p},$
@@
    Fix $z>0$ and note $M_{\tau,\lambda}=A_{\tau,\lambda}\wedge S_{\tau,\lambda}+\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert$. By 5.3(d), the source condition on $f^{\dagger}$ and (5.4) we have
@@
    $\leqslant\inf_{\rho>0}\Big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+\delta^{2}\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)Z_{i}^{2}\Big{)}.$ (9.6)
    Choosing $\rho\thicksim(R\delta^{-1})^{4p/(4\mu p+2p+1)}$ gives the right order of the first summand. For the second summand, we obtain similarly to (5.6)
    $\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)^{2}\leqslant C_{A}^{4}\sum_{i=1}^{D}\rho^{2}i^{-4p}\wedge 1\leqslant C_{A}^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}.$
@@
    $\operatorname{{\mathbb{P}}}\Big{(}\textstyle\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)(Z_{i}^{2}-1)\geqslant 2C_{A}^{2}\sqrt{\tfrac{4p}{4p-1}}\rho^{1/(4p)}\sqrt{z}+2z\Big{)}\leqslant e^{-z}.$
    From $2C_{A}^{2}\sqrt{\tfrac{4p}{4p-1}}\rho^{1/(4p)}\sqrt{z}+2z\leqslant C_{A}^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}+3z$ we deduce with (5.6)
    $\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)Z_{i}^{2}\leqslant\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)+C_{A}^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}+3z\leqslant(1\vee C_{A}^{4})\tfrac{4p}{2p-1}\rho^{1/(2p)}+3z$
@@
    Fix $z\geqslant 1$ and put $\Delta:=\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert$ for short. Using (6.1), the identity $R_{\tau}^{2}=\kappa$ and the assumption on $\kappa$, we bound
    $\Delta$ $\leqslant\big{\lvert}\lVert\xi\rVert^{2}-\delta^{2}D\big{\rvert}+\lvert\kappa-\delta^{2}D\rvert+2\lvert\langle{\xi},{\mathsf{r_{\tau,<}}g}\rangle\rvert$
@@
[…]
@@
    and the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, we obtain
    $\lVert\mathsf{\exp(-\lambda_{D}^{-2}x)}g\rVert\leqslant\big{(}(\mu+1/2)e^{-1}\big{)}^{\mu+1/2}\lambda_{D}^{2\mu+1}R\lesssim_{\mu,C_{A}}RD^{-2\mu p-p}.$
@@
    From the assumption $D\gtrsim(R\delta^{-1})^{2/(4\mu p+2p+1)}$ we thus derive (9.10) again. ∎
@@
Fix $z\geqslant 1$. By 7.8(a), the inequality
@@
holds true with probability at least $1-e^{-z}$. We control the summands using 7.8(c). The following bounds hold simultaneously with probability at least $1-5e^{-z}$:
@@
By Young’s inequality with $\tilde{p}\coloneqq(2p+1)/(2p)$ and $\tilde{q}\coloneqq 2p+1$, we deduce
@@
Proposition 5.14 implies
@@
Noting that $\lVert\mathsf{r_{\tau,<}}g\rVert\vee\lVert A(\widehat{f}_{\tau}-f)\rVert\leqslant R_{\tau}+\lVert\xi\rVert\leqslant(c+1)\delta$ and $S_{\tau,\lambda}\leqslant\delta^{2}$ under (DN), we obtain the claim by plugging in the bound on $\lvert r_{\tau}^{\prime}(0)\rvert$. ∎
@@ Acknowledgements
Financial support by the Deutsche Forschungsgemeinschaft (DFG) through the research unit FOR 5381 Mathematical Statistics in the Information Age – Statistical Efficiency and Computational Tractability is gratefully acknowledged.

### v2

@@ 9.4 Proofs for Section 5
We prove the claim analogously to Johnstone [15, Proposition 4.23]. Without loss of generality, assume that $v_{1},\dots,v_{D}$ are the first $D$ canonical unit vectors. Let $\Theta^{m}(\eta)$, $m=1,\dots,D$, for $\eta\geqslant 0$ denote the set of $f\in{\mathbb{R}}^{P}$ that satisfy $\lvert f_{i}\rvert\leqslant\eta$ for $i=1,\dots,m$ and $f_{i}=0$ for $i=m+1,\dots,P$ in case of $P>m$. Define the loss
@@
Hence, for $\Theta^{m}(\delta m^{p})$ to be contained in the class of signals $f$ satisfying $\lVert f\rVert_{\mu}\leqslant R$, it suffices that $c_{A}^{-4\mu}\delta^{2}m^{4\mu p+2p+1}\leqslant R^{2}$. Let $m_{1}$ be the largest integer such that $c_{A}^{-4\mu}m_{1}^{4\mu p+2p+1}\leqslant R^{2}/\delta^{2}$ and $m_{0}\in{\mathbb{R}}$ the solution to $c_{A}^{-4\mu}m_{0}^{4\mu p+2p+1}=R^{2}/\delta^{2}$. If $m_{0}\geqslant 4$, then $m_{1}/m_{0}\geqslant 1/2$, which can be shown by contradiction. For $\delta/R\leqslant\tilde{c}_{\mu,p,c_{A}}$, we conclude
@@ 9.5 Proofs for Section 7
    Consider the event of Lemma 6.4, which holds with probability at least $1-e^{-z}$. If $\lvert r_{\tau}^{\prime}(0)\rvert<\lambda_{D}^{-2}$, then by the first implication of Lemma 6.4
@@
    and the assertion follows from Corollary 7.3. If $\lvert r_{\tau}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, then we use the rough bound
@@
    which is due to the prediction error bound from Proposition 5.2. Moreover, we then have by PSD($p$,$c_{A}$,$C_{A}$) and the second implication of Lemma 6.4
    $\lambda_{D}^{-2}\lesssim_{c_{A}}D^{2p}\lesssim_{p,c_{A}}(\delta^{-2}S_{\tau,\lambda})^{2p}+(\log d)^{2p}+z^{2p},$
@@
    Fix $z>0$ and note $M_{\tau,\lambda}=A_{\tau,\lambda}\wedge S_{\tau,\lambda}+\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert$. In the case $\tau>0$, we have by 5.3(c), the source condition on $f$ and (5.4)
@@
    $\leqslant\inf_{\rho>0}\Big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+\delta^{2}\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)Z_{i}^{2}\Big{)},$ (9.7)
    where $Z_{i}\coloneqq\langle{Z},{u_{i}}\rangle$, $i=1,\dots,D$. For $\tau=0$, the bound (9.7) continues to hold due to $A_{\tau,\lambda}\wedge S_{\tau,\lambda}=0$. Choosing $\rho\thicksim(R\delta^{-1})^{4p/(4\mu p+2p+1)}$ gives the right order of the first summand. For the second summand, we obtain similarly to (5.6)
    $\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)^{2}\leqslant(C_{A}\vee 1)^{4}\sum_{i=1}^{D}(\rho^{2}i^{-4p}\wedge 1)\leqslant(C_{A}\vee 1)^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}.$
@@
    $\operatorname{{\mathbb{P}}}\Big{(}\textstyle\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)(Z_{i}^{2}-1)\geqslant 2(C_{A}\vee 1)^{2}\sqrt{\tfrac{4p}{4p-1}}\rho^{1/(4p)}\sqrt{z}+2z\Big{)}\leqslant e^{-z}.$
    From $2(C_{A}\vee 1)^{2}\sqrt{\tfrac{4p}{4p-1}}\rho^{1/(4p)}\sqrt{z}+2z\leqslant(C_{A}\vee 1)^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}+3z$ we deduce with (5.6)
    $\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)Z_{i}^{2}$ $\leqslant\sum_{i=1}^{D}(\rho\lambda_{i}^{2}\wedge 1)+(C_{A}\vee 1)^{4}\tfrac{4p}{4p-1}\rho^{1/(2p)}+3z$
    $\leqslant(C_{A}\vee 1)^{4}\tfrac{4p}{2p-1}\rho^{1/(2p)}+3z$
@@
    Fix $z\geqslant 1$ and put $\Delta:=\lvert A_{\tau,\lambda}-S_{\tau,\lambda}\rvert$ for short. In the case $\kappa\leqslant\lVert Y\rVert^{2}$, we have $R_{\tau}^{2}=\kappa$ such that by (6.1)
    $\Delta=\lvert\kappa-\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{\tau,<}}g}\rangle\rvert.$
    Otherwise, if $\kappa>\lVert Y\rVert^{2}$, we stop at $\tau=0$ and
    $0\leqslant\Delta=A_{0,\lambda}=\lVert g\rVert^{2}$ $=\lVert Y\rVert^{2}+\lVert\xi\rVert^{2}-2\langle{\xi},{Y}\rangle$
    $=\lVert Y\rVert^{2}-\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{0,<}}g}\rangle<\kappa-\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{0,<}}g}\rangle.$
    Hence, using the assumption on $\kappa$, we can in both cases bound
    $\Delta$ $\leqslant\lvert\kappa-\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{\tau,<}}g}\rangle\rvert$
    $\leqslant\big{\lvert}\lVert\xi\rVert^{2}-\delta^{2}D\big{\rvert}+\lvert\kappa-\delta^{2}D\rvert+2\lvert\langle{\xi},{\mathsf{r_{\tau,<}}g}\rangle\rvert$
@@
    In the case $\lvert r_{\tau}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, we have by Lemma 4.7
    $S_{\tau,\lambda}\geqslant\lVert\mathsf{(1-{\exp(-\lambda_{D}^{-2}x)})^{1/2}}\xi\rVert^{2}\geqslant(1-e^{-1})\lVert\xi\rVert^{2}$
    and bound directly
@@
    and the source condition $\lVert f\rVert_{\mu}\leqslant R$, we obtain
    $\lVert\mathsf{\exp(-\lambda_{D}^{-2}x)}g\rVert$ $=\Big{(}\sum_{i=1}^{D}\big{(}\exp(-2\lambda_{D}^{-2}\lambda_{i}^{2})\lambda_{i}^{4\mu+2}\big{)}\lambda_{i}^{-4\mu}f_{i}^{2}\Big{)}^{1/2}$
    $\leqslant\big{(}(\mu+1/2)e^{-1}\big{)}^{\mu+1/2}\lambda_{D}^{2\mu+1}R\lesssim_{\mu,C_{A}}RD^{-2\mu p-p}.$
@@
    From the assumption $D\gtrsim(R\delta^{-1})^{2/(4\mu p+2p+1)}$ we thus derive (9.11) again on the same event. ∎
@@
Fix $z\geqslant 1$ and recall $M_{\tau,\lambda}\coloneqq A_{\tau,\lambda}\vee S_{\tau,\lambda}$. By 7.8(a), the inequality
@@
holds true with probability at least $1-e^{-z}$. We control the summands using 7.8(c). The following bounds hold simultaneously with probability at least $1-5e^{-z}$, where $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ is given in (3.5):
@@
Finally, by Young’s inequality with $\tilde{p}\coloneqq(2p+1)/(2p)$ and $\tilde{q}\coloneqq 2p+1$, we deduce
@@
In the case $c\delta\leqslant\lVert Y\rVert$, we have $R_{\tau}=c\delta$ and Proposition 5.14 implies
@@
Noting that $\lVert\mathsf{r_{\tau,<}}g\rVert\vee\lVert A(\widehat{f}_{\tau}-f)\rVert\leqslant R_{\tau}+\lVert\xi\rVert\leqslant(c+1)\delta$ and $S_{\tau,\lambda}\leqslant\delta^{2}$ under (DN), we obtain the claim by plugging in the bound on $\lvert r_{\tau}^{\prime}(0)\rvert$. Otherwise, if $c\delta>\lVert Y\rVert$, we stop at $\tau=0$ and obtain similar to (7.3)
$\lVert\widehat{f}_{\tau}-f^{\dagger}\rVert^{2}=\lVert f^{\dagger}\rVert^{2}\leqslant\lVert g\rVert^{4\mu/(2\mu+1)}R^{2/(2\mu+1)}.$
Plugging in $\lVert g\rVert\leqslant\lVert Y\rVert+\lVert\xi\rVert<(c+1)\delta$ yields the claim. ∎
@@ Acknowledgements
We are grateful for comments by Tim Jahn that helped to improve the paper. Financial support by the Deutsche Forschungsgemeinschaft (DFG) through the research unit FOR 5381 Mathematical Statistics in the Information Age – Statistical Efficiency and Computational Tractability is gratefully acknowledged.